OpenStudy (fairy733189):

The table below shows the sale, in dollars, at Jacob's store over a period of five months: month: 1 2 3 4 5 sale: 1000 1050 1102.50 1,157.63 1215.51 Did the number of people at Jacob's store increase linearly or exponentially? (4 points) Linearly, because the table shows a constant percentage increase in sales per month Exponentially, because the table shows a constant percentage increase in sales per month Linearly, because the table shows that sales increase by an equal factor for an equal increase in months Exponentially because the table shows an equal increase in sales for an equal increase in months

OpenStudy (fairy733189):

@vishweshshrimali5

OpenStudy (vishweshshrimali5):

To see whether the increase is linear or exponential... Do one thing.. calculate the slope between two consecutive points... Slope = Change in sale/ Change in number of months

OpenStudy (vishweshshrimali5):

Can you do this?

OpenStudy (fairy733189):

Does that mean I have to find the percentage increase?

OpenStudy (vishweshshrimali5):

No .. percentage is not required... Just relative change

OpenStudy (fairy733189):

Then no I don't know how to do this

OpenStudy (vishweshshrimali5):

Okay let me show you for one set of points... Month: 1 2 Sale: 1000 1050 Now, change in sale = 1050 - 1000 = 50 Change in number of months = 2 - 1 = 5 Slope = 50 / 5 = 10

OpenStudy (vishweshshrimali5):

Can you do this for rest of the points?

OpenStudy (fairy733189):

Sure but it'll take a few minutes

OpenStudy (vishweshshrimali5):

Find the slope between, ( 2 and 3), (3 and 4), (4 and 5)

OpenStudy (fairy733189):

Wait I just have a question... Why is the change in the number of months 5?

OpenStudy (vishweshshrimali5):

Hmm? Between month 4 and month 5 you mean?

OpenStudy (fairy733189):

no between 1 and 2

OpenStudy (vishweshshrimali5):

Ohh sorry that's my mistake :)

OpenStudy (vishweshshrimali5):

That should be change in number of months = 1 Slope = 50 / 1 = 50

OpenStudy (fairy733189):

ok thanks

OpenStudy (fairy733189):

now i'll just continue doing the rest

OpenStudy (vishweshshrimali5):

Sure

OpenStudy (fairy733189):

ok so for months 2 and 3 the slope = 52.5 3 and 4 the slope = 55.3 4 and 5 the slope = 57.88

OpenStudy (vishweshshrimali5):

Great! So, you see that the relative change (and thus percentage change) is not constant so the change is exponential

OpenStudy (fairy733189):

One more question before moving on. It's something that always confuses me. What is the difference between linear, or constant, and exponential?

OpenStudy (fairy733189):

Sorry if that's a dumb question. I'm just not that good at math.

OpenStudy (vishweshshrimali5):

Actually, linear change means that the changes are increasing is with equal amount... Constant means that changes are same Exponential means that difference between changes is increasing in a non-equal amount ...

OpenStudy (fairy733189):

Oh. Thanks for explaining. Can I just have like 2 minutes to quickly write this down?

OpenStudy (vishweshshrimali5):

Sure! Take your time :)

OpenStudy (fairy733189):

ok thanks i'm done!

OpenStudy (fairy733189):

@vishweshshrimali5

OpenStudy (vishweshshrimali5):

So, @Fairy733189 you want to ask anything?

OpenStudy (fairy733189):

no I actually figure it out and I understand it now but can you help me on another problem